# In-context Learning with Retrieved Demonstrations for Language Models: A Survey

### 13 Feb 2024 v3

@@ 1 Introduction
Early ICL implementations use a fixed set of demonstrations for each target task. These demonstrations could be hand-crafted by human (Hendrycks et al., 2021; Wei et al., 2022; Kazemi et al., 2023b), randomly chosen from training data (Brown et al., 2020; Lewkowycz et al., 2022), or chosen based on metrics such as complexity or information content (Fu et al., 2022; Hongjin et al., 2022; Li and Qiu, 2023a; Wang et al., 2023b). The effectiveness of such demonstrations is influenced by factors such as the quality, quantity, and ordering of the demonstrations. Importantly, the demonstrations remain context-insensitive (i.e. the same demonstrations are used regardless of the query) which could hinder unlocking the true potential of the LLMs.
@@
Table 1: Comparison with Related Work. Abbreviation for Evaluation Tasks: CodeGen (code generation), SA (sentiment analysis), Table2Text (Table to Text generation), QA (question answering), SP (semantic parsing), DST (Dialogue State Tracking), D2T (Data-to-Text), Summ ( Summarization), CSR ( commonsense reasoning), RC (reading comprehension), NLI (natural language inference), CR (Coreference Resolution), MathR (mathematical reasoning), PD (paraphrase detection), TQA (Table Question Answering), TC (Topic Classification), StoryGen (Story Generation).
@@ 2.2 What Makes for Good Demonstrations?
LLMs generally benefit from more demonstrations, but as the number of demonstrations increases the rate of improvement typically decreases (Brown et al., 2020; Ye et al., 2023b; Min et al., 2022b). Generation tasks have been shown to benefit from an increased number of demonstrations more than classification tasks (Li et al., 2023). Toward increasing the number of demonstrations, one barrier is the maximum context size of the LLM. While the size of the context has been increasing over time with newer LLMs, it may still be problematic for datasets with long input texts or classification datasets with many classes.
@@ 3 In-context Learning with Demonstration Retrieval
Traditionally, the same set of few-shot demonstrations is used on all queries, which can be suboptimal especially when there are high variations among the queries. An alternative is to retrieve few-shot demonstrations that are tailored to the current query. Previous work has shown that demonstration retrieval leads to substantial improvements in the task metrics, compared to manually curated or randomly selected demonstrations (Luo et al., 2023; Ye et al., 2023a). Furthermore, LLMs have been shown to become less sensitive to the factors such as demonstration ordering (Section 2.2) when retrieved demonstrations are used (Li et al., 2023).
@@ 3.1 Retrieval Objectives: Similarity and Diversity
Various retrieval objectives for selecting and tailoring in-context examples for LLMs have been explored (Luo et al., 2023; Rubin et al., 2022; Ye et al., 2023a; Dalvi et al., 2022; Cheng et al., 2023; Li et al., 2023). There are two primary retrieval objectives for selecting demonstrations: similarity and diversity. Similarity involves selecting demonstrations most akin to the query and can be based on language similarity (term matching or semantic matching), structural aspects (sentence structure, reasoning structure, etc.), or other criteria. Most studies focus on language similarity, with fewer addressing structural similarity, often due to the challenges in extracting a query’s structure in many tasks (Levy et al., 2022). Beyond similarity, some work has found that the diversity of demonstrations is important. The motivations for diversity include avoiding repetitive demonstrations (Zhang et al., 2022b), bringing different perspectives (Yu et al., 2023), and maximizing the demonstrations’ coverage of the test query, in terms of covering either its words or syntactic structures (Levy et al., 2022). Measuring the diversity of multiple demonstrations is a major technical challenge.  Ye et al. (2023a) applied determinantal point processes (DPP) a probabilistic model to measure the negative interaction (Kulesza et al., 2012), to measure the diversity.  Levy et al. (2022) found that diversity and coverage are important when the model is unfamiliar with the output symbols space. It is noteworthy that researchers have found that ICL benefits more from demonstrations with higher complexity in some scenarios (Fu et al., 2022), where they define the complexity in terms of the query length or reasoning steps. However,  Fu et al. (2022) employed heuristic rules to define complexity and pre-selected demonstrations accordingly. Their research revealed that using a similarity-based retriever led to improved performance in a specific mathematical reasoning task. This might indicate that combining similarity and complexity considerations could be a promising strategy for enhancing the approach to reasoning tasks.
@@ 3.3 Retrieval Corpus
The previous scenario has one individual retrieval corpus for different tasks. Assuming that we want to test model performance on two tasks, then in the in-domain setting, there will be two retrieval corpora separately. Furthermore, the in-domain setting assumes that the model has knowledge about which task the test question belongs to such that when it comes to the retrieval phase, it knows which corpus to select the demonstrations from. However, this assumption does not hold in several real-world applications of LLMs. In the mix-domain setting (Wang et al., 2023a; Li et al., 2023), the retrieval corpus is constructed from the combination of all tasks. At the inference time, given a question, the retriever will retrieve demonstrations from this mixed corpus; the demonstrations can come from the same domain as the test question or from other tasks.
@@ 4 Off-the-shelf Demonstration Retrievers
In this approach, queries and documents are encoded to the same dense embedding space using an off-the-shelf sentence embedding model, and then similarity scores (e.g. cosine similarity) are calculated to rank the most relevant documents for each query. A rich collection of sentence embedding methodologies exists in the literature. Sentence-BERT (SBERT) (Reimers and Gurevych, 2019a) is a modification of the pretrained BERT network that uses siamese and triplet network structures to derive semantically meaningful sentence embeddings. The effectiveness of SBERT embeddings for demonstration retrieval has been investigated in several works (Rubin et al., 2022; Li and Qiu, 2023b; Wang et al., 2023a), and the results show that retrieving demonstrations based on SBERT embeddings often provides a boost in performance compared to zero-shot or random few-shot selection. In the KATE method (Liu et al., 2022), the authors studied using vanilla RoBERTa (Liu et al., 2019) and finetuned RoBERTa on NLI (Bowman et al., 2015b) and STS-B (Cer et al., 2017) datasets for selecting good demonstrations, and found that the finetuned version on task-related datasets offered further empirical gains. Note that here, the demonstration retriever is not trained for ICL demonstration retrieval based on task-specific data (a topic which we will discuss in Section 5); instead, the retriever is finetuned related tasks to provide a better notion of similarity for the task at hand. So we still categorize it as an off-the-shelf retriever. Shi et al. (2022) extends the use case to cross-lingual few-shot retrieval in the Text to-SQL semantic parsing task, and they use mSBERT (Reimers and Gurevych, 2019b), mUSE (Yang et al., 2019) and mT5 (Xue et al., 2020) as the baseline models for comparison. Other widely used baseline models for demonstration retrieval include $E5_{\text{base}}$ (Wang et al., 2022b), SimCSE (Gao et al., 2021). Instead of relying on “word matches” as in BM25, these sentence embedding similarity approaches can better capture semantic similarity (for example, synonyms, and related topics), however computationally they might be more expensive.
@@ 5.1 Collecting Training Data for Demonstration Retriever
There are different strategies for choosing the scoring LLM. Ideally, one uses the inference LLM itself as the scorer in order to perfectly reflect its preferences (Li et al., 2023; Shi et al., 2022). However, training retrievers requires large amounts of labeled data, and it may be expensive use very large models for labeling. Consequently, for scoring one may gravitate towards utilizing smaller models, especially those within the same model family as the inference LLM (Luo et al., 2023; Cheng et al., 2023; Rubin et al., 2022).
@@ 5.2 Training Objectives
The list-wise ranking approach looks at a list of candidate documents for a given query and tries to capture the correct ordering for it. Li et al. (2023) proposed to inject the ranking signals into the retriever using an approach inspired by LambdaRank (Burges, 2010). More formally, given each query $q$, they first rank all $l$ candidate documents according to their relevant scores $S=\{s(d_{i})\}_{i=1}^{l}$, according to which the associated ranking =$\mathcal{R}=\{r(d_{i})\}_{i=1}^{l}$ is computed. Then the loss function is defined as follows:
@@
Here $\alpha$ is a constant that controls the relative importance of the two losses. They claimed that with the multi-objective function, both the absolute scores and supervised signals are taken into consideration. Li et al. (2023) trains a universal retriever with both list-wise ranking loss and the InfoCNE loss.
@@
Regarding training strategies, most research efforts have centered on fine-tuning a single retriever. Wang et al. (2023a) and Li et al. (2023) instead proposed to iterate the retriever model multiple times. More specifically, the retriever trained in iteration $i$ will be employed to retrieve a new set of candidates for the subsequent iteration $i+1$. Such an iterative training approach allows progressively improving retriever quality by mining better positive and hard negative examples at each iteration.
@@ 5.3 Summary
Here, we summarize the advantages and disadvantages of various retriever models. The off-the-shelf retrievers are easy to use without any downstream task finetuning and typically demonstrate stronger performance than random demonstrations. One exception is in commonsense reasoning tasks where Zhang et al. (2022b) and Ye et al. (2023a) found that for these tasks, random demonstrations are consistently better than retrieval-based method. Cheng et al. (2023) also show that retrieved demonstrations harm commonsense reasoning and coreference resolution tasks. Among the three categories of off-the-shelf retrievers, sparse retrievers such as BM25 are more index-efficient. This feature becomes particularly valuable when dealing with large volumes of demonstrations and limited hardware memory, making BM25 a preferable choice under such circumstances. In contrast, sentence-embedding similarity-based methods and dual-encoder-based retrieval systems, which are trained on language tasks, excel in capturing more semantically focused retrieval. Regarding performance,  Luo et al. (2023) compared BM25 with dual encoder (GTR) across 5 tasks, and they found that the average performance of these two is very similar (within 0.5% difference), and BM25 outperformed the dual encoder in some tasks and vice versa. In another study, Ye et al. (2023a) observed a similar trend highlighting that no single retriever consistently outperforms others across different tasks. Both Rubin et al. (2022) and Li et al. (2023) found that BM25 is better than SBERT on semantic parsing tasks, while  Li et al. (2023) found that SBERT is better than BM25 on sentiment analysis tasks. Nevertheless, retrievers that are fine-tuned demonstrate superior performance compared to their off-the-shelf counterparts. The main drawback of fine-tuned retrievers lies in the high cost of obtaining training data. Additionally, the common practice of employing task-specific retrievers complicates the system and limits its generalizability.  Li et al. (2023) proposed to train a universal retriever that shows stronger performance than task-specific demonstration retriever (e.g. EPR (Rubin et al., 2022)) on most of the tasks.

### 12 Mar 2024 v4

@@ 1 Introduction
Early ICL implementations use a fixed set of demonstrations for each target task. These demonstrations could be hand-crafted by human (Hendrycks et al., 2021; Wei et al., 2022; Kazemi et al., 2023b), randomly chosen from training data (Brown et al., 2020; Lewkowycz et al., 2022). Beyond random selection, there are more advanced selection processes based on metrics such as complexity (Fu et al., 2022), diversity (Li and Qiu, 2023a), difficulty (Drozdov et al., 2023), concept learning (Wang et al., 2023b) and perplexity (Gonen et al., 2023). Importantly, the demonstrations remain context-insensitive (i.e. the same demonstrations are used regardless of the query) which could hinder unlocking the true potential of the LLMs. The effectiveness of such demonstrations is influenced by factors such as the quality, quantity, and ordering of the demonstrations.
@@
Dynamic Least-to-Most 2022
DODEX
Tree Structure Similarity
In Domain
SP
Diversity and Lexical Overalp
Iterative
@@
R-BM25 2023
XGLM-7.5G
BM25, BM25 and Term Recall based rerank In Domain
In Domain
MT
Term-based Similarity
One-hoc
ICL-MC 2023
GPTNeo-2.7B, XGLM-2.9G, BLOOM-3B
BM25
In Domain Cross Domain
MT
Term-based Similarity
One-hoc
ICL-ML 2023
OPT-13/175B, LLaMA-7/70B
SBERT
In Domain
Intent Classification
Sentence Embeddings Similarity
One-hoc
@@
Table 1: Comparison with Related Work. Abbreviation for Evaluation Tasks: CodeGen (code generation), SA (sentiment analysis), Table2Text (Table to Text generation), QA (question answering), SP (semantic parsing), DST (Dialogue State Tracking), D2T (Data-to-Text), Summ ( Summarization), CSR ( commonsense reasoning), RC (reading comprehension), NLI (natural language inference), CR (Coreference Resolution), MathR (mathematical reasoning), PD (paraphrase detection), TQA (Table Question Answering), TC (Topic Classification), StoryGen (Story Generation), MT (Machine Translation)
@@ 2.2 What Makes for Good Demonstrations?
LLMs generally benefit from more demonstrations, but as the number of demonstrations increases the rate of improvement typically decreases (Brown et al., 2020; Ye et al., 2023b; Min et al., 2022b). Generation tasks have been shown to benefit from an increased number of demonstrations more than classification tasks (Li et al., 2023b). Toward increasing the number of demonstrations, one barrier is the maximum context size of the LLM. While the size of the context has been increasing over time with newer LLMs, it may still be problematic for datasets with long input texts or classification datasets with many classes.
@@ 3 In-context Learning with Demonstration Retrieval
Traditionally, the same set of few-shot demonstrations is used on all queries, which can be suboptimal especially when there are high variations among the queries. An alternative is to retrieve few-shot demonstrations that are tailored to the current query. Previous work has shown that demonstration retrieval leads to substantial improvements in the task metrics, compared to manually curated or randomly selected demonstrations (Luo et al., 2023; Ye et al., 2023a). Furthermore, LLMs have been shown to become less sensitive to the factors such as demonstration ordering (Section 2.2) when retrieved demonstrations are used (Li et al., 2023b).
@@ 3.1 Retrieval Objectives: Similarity and Diversity
Various retrieval objectives for selecting and tailoring in-context examples for LLMs have been explored (Luo et al., 2023; Rubin et al., 2022; Ye et al., 2023a; Dalvi et al., 2022; Cheng et al., 2023; Li et al., 2023b). There are two primary retrieval objectives for selecting demonstrations: similarity and diversity. Similarity involves selecting demonstrations most akin to the query and can be based on language similarity (term matching or semantic matching), structural aspects (sentence structure, reasoning structure, etc.), or other criteria. Most studies focus on language similarity, with fewer addressing structural similarity, often due to the challenges in extracting a query’s structure in many tasks (Levy et al., 2022). Beyond similarity, some work has found that the diversity of demonstrations is important. The motivations for diversity include avoiding repetitive demonstrations (Zhang et al., 2022b), bringing different perspectives (Yu et al., 2023), and maximizing the demonstrations’ coverage of the test query, in terms of covering either its words or syntactic structures (Levy et al., 2022). Measuring the diversity of multiple demonstrations is a major technical challenge.  Ye et al. (2023a) applied determinantal point processes (DPP) a probabilistic model to measure the negative interaction (Kulesza et al., 2012), to measure the diversity.  Levy et al. (2022) found that diversity and coverage are important when the model is unfamiliar with the output symbols space. It is noteworthy that researchers have found that ICL benefits more from demonstrations with higher complexity in some scenarios (Fu et al., 2022), where they define the complexity in terms of the query length or reasoning steps. However,  Fu et al. (2022) employed heuristic rules to define complexity and pre-selected demonstrations accordingly. Their research revealed that using a similarity-based retriever led to improved performance in a specific mathematical reasoning task. This might indicate that combining similarity and complexity considerations could be a promising strategy for enhancing the approach to reasoning tasks.
@@ 3.3 Retrieval Corpus
The previous scenario has one individual retrieval corpus for different tasks. Assuming that we want to test model performance on two tasks, then in the in-domain setting, there will be two retrieval corpora separately. Furthermore, the in-domain setting assumes that the model has knowledge about which task the test question belongs to such that when it comes to the retrieval phase, it knows which corpus to select the demonstrations from. However, this assumption does not hold in several real-world applications of LLMs. In the mix-domain setting (Wang et al., 2023a; Li et al., 2023b), the retrieval corpus is constructed from the combination of all tasks. At the inference time, given a question, the retriever will retrieve demonstrations from this mixed corpus; the demonstrations can come from the same domain as the test question or from other tasks.
@@ 4 Off-the-shelf Demonstration Retrievers
In this approach, queries and documents are encoded to the same dense embedding space using an off-the-shelf sentence embedding model, and then similarity scores (e.g. cosine similarity) are calculated to rank the most relevant documents for each query. A rich collection of sentence embedding methodologies exists in the literature. Sentence-BERT (SBERT) (Reimers and Gurevych, 2019a) is a modification of the pretrained BERT network that uses siamese and triplet network structures to derive semantically meaningful sentence embeddings. The effectiveness of SBERT embeddings for demonstration retrieval has been investigated in several works (Rubin et al., 2022; Li and Qiu, 2023b; Wang et al., 2023a), and the results show that retrieving demonstrations based on SBERT embeddings often provides a boost in performance compared to zero-shot or random few-shot selection. In the KATE method (Liu et al., 2022), the authors studied using vanilla RoBERTa (Liu et al., 2019) and finetuned RoBERTa on NLI (Bowman et al., 2015b) and STS-B (Cer et al., 2017) datasets for selecting good demonstrations, and found that the finetuned version on task-related datasets offered further empirical gains. Note that here, the demonstration retriever is not trained for ICL demonstration retrieval based on task-specific data (a topic which we will discuss in Section 5); instead, the retriever is finetuned related tasks to provide a better notion of similarity for the task at hand. So we still categorize it as an off-the-shelf retriever. Shi et al. (2022) extends the use case to cross-lingual few-shot retrieval in the Text to-SQL semantic parsing task, and they use mSBERT (Reimers and Gurevych, 2019b), mUSE (Yang et al., 2019) and mT5 (Xue et al., 2020) as the baseline models for comparison. Other widely used baseline models for demonstration retrieval include $E5_{\text{base}}$ (Wang et al., 2022b), SimCSE (Gao et al., 2021b). Instead of relying on “word matches” as in BM25, these sentence embedding similarity approaches can better capture semantic similarity (for example, synonyms, and related topics), however computationally they might be more expensive.
@@ 5.1 Collecting Training Data for Demonstration Retriever
There are different strategies for choosing the scoring LLM. Ideally, one uses the inference LLM itself as the scorer in order to perfectly reflect its preferences (Li et al., 2023b; Shi et al., 2022). However, training retrievers requires large amounts of labeled data, and it may be expensive use very large models for labeling. Consequently, for scoring one may gravitate towards utilizing smaller models, especially those within the same model family as the inference LLM (Luo et al., 2023; Cheng et al., 2023; Rubin et al., 2022).
@@ 5.2 Training Objectives
The list-wise ranking approach looks at a list of candidate documents for a given query and tries to capture the correct ordering for it. Li et al. (2023b) proposed to inject the ranking signals into the retriever using an approach inspired by LambdaRank (Burges, 2010). More formally, given each query $q$, they first rank all $l$ candidate documents according to their relevant scores $S=\{s(d_{i})\}_{i=1}^{l}$, according to which the associated ranking =$\mathcal{R}=\{r(d_{i})\}_{i=1}^{l}$ is computed. Then the loss function is defined as follows:
@@
Here $\alpha$ is a constant that controls the relative importance of the two losses. They claimed that with the multi-objective function, both the absolute scores and supervised signals are taken into consideration. Li et al. (2023b) trains a universal retriever with both list-wise ranking loss and the InfoCNE loss.
@@
Regarding training strategies, most research efforts have centered on fine-tuning a single retriever. Wang et al. (2023a) and Li et al. (2023b) instead proposed to iterate the retriever model multiple times. More specifically, the retriever trained in iteration $i$ will be employed to retrieve a new set of candidates for the subsequent iteration $i+1$. Such an iterative training approach allows progressively improving retriever quality by mining better positive and hard negative examples at each iteration.
@@ 5.3 Summary
Here, we summarize the advantages and disadvantages of various retriever models. The off-the-shelf retrievers are easy to use without any downstream task finetuning and typically demonstrate stronger performance than random demonstrations. One exception is in commonsense reasoning tasks where Zhang et al. (2022b) and Ye et al. (2023a) found that for these tasks, random demonstrations are consistently better than retrieval-based method. Cheng et al. (2023) also show that retrieved demonstrations harm commonsense reasoning and coreference resolution tasks. Among the three categories of off-the-shelf retrievers, sparse retrievers such as BM25 are more index-efficient. This feature becomes particularly valuable when dealing with large volumes of demonstrations and limited hardware memory, making BM25 a preferable choice under such circumstances. In contrast, sentence-embedding similarity-based methods and dual-encoder-based retrieval systems, which are trained on language tasks, excel in capturing more semantically focused retrieval. Regarding performance,  Luo et al. (2023) compared BM25 with dual encoder (GTR) across 5 tasks, and they found that the average performance of these two is very similar (within 0.5% difference), and BM25 outperformed the dual encoder in some tasks and vice versa. In another study, Ye et al. (2023a) observed a similar trend highlighting that no single retriever consistently outperforms others across different tasks. Both Rubin et al. (2022) and Li et al. (2023b) found that BM25 is better than SBERT on semantic parsing tasks, while  Li et al. (2023b) found that SBERT is better than BM25 on sentiment analysis tasks. Nevertheless, retrievers that are fine-tuned demonstrate superior performance compared to their off-the-shelf counterparts. The main drawback of fine-tuned retrievers lies in the high cost of obtaining training data. Additionally, the common practice of employing task-specific retrievers complicates the system and limits its generalizability.  Li et al. (2023b) proposed to train a universal retriever that shows stronger performance than task-specific demonstration retriever (e.g. EPR (Rubin et al., 2022)) on most of the tasks.
@@ 7 Discussion of Future Direction
Fine-tuning For Few-shot Learning
Most of the existing research utilizes a frozen LLM as the few-shot learner in the RetICL framework. As demonstrated by Gao et al. (2021a); Izacard et al. (2023), fine-tuning LLMs on a few-shot learning task can be a promising approach to enhancing LLM performance during inference. An intriguing avenue could involve adapting the fine-tuning strategy from open-domain question answering models to RetICL (Lewis et al., 2020; Guu et al., 2020).
RetICL for Vision and Language Models
Vision and language models (VL) have demonstrated proficiency as few-shot learners (Alayrac et al., 2022; Awadalla et al., 2023; Li et al., 2023a), and researchers have integrated retrieval augmented generation methods with VL models (Luo et al., 2021; Gao et al., 2022; Yasunaga et al., 2023). In lieu of employing random demonstrations, Yang et al. (2024) leverages a retriever to select demonstrations for image captioning tasks. Additionally, Peng et al. (2023) propose an ICD-LM to generate demonstrations for a given query, with each demonstration being represented by a token. The effectiveness of this method is evidenced by its performance on image captioning and vision question answering tasks. Despite the increasing significance of vision and language generation models in real-world applications, research on RetICL for VL models remains relatively underexplored.
@@ 8 Conclusion
Acknowledge
We would like to thank Siva Reddy, Suzanna Sia, Andrew Drozdov and Yang Xu for recommending additional references for the initial draft. Special thanks are extended to Andrew Drozdov for suggesting the direction of future work on fine-tuning for few-shot learning, and to Xu Yang for suggesting RetICL for VL models.